A gaseous fuel has the following composition by volume c2h4=20%,c4h10=30%,co=15%,o2=2%.calculate the volume of air supplied per m 3 of fuel


1
Expert's answer
2021-02-18T07:31:19-0500

Suppose the conditions are normal: T = 20 OC = 293 OK, P = 101.325 kPa. Also, the air contain 21% O2. Also, 20%+30%+15+2% = 67%. Let's say the rest 33% is an inflamable gas, for example N2.


ngas = "\frac{PV}{RT}" = "\frac{101.325 \times 1000}{8.314 \times 293}" = 40.90 (mol)


n(C2H4) = 40.90 "\times" 0.2 = 8.18 (mol)

n(C4H10) = 40.90 "\times" 0.3 = 12.27 (mol)

n(CO) = 40.90 "\times" 0.15 = 6.13 (mol)

n(O2) = 40.90"\times"0.02 = 0.82 (mol)


C2H4 + 3O2 = 2CO2 + 2H2O

C4H10 + 6"\frac{1}{2}"O2 = 4CO2 + 5H2O

CO + "\frac{1}{2}"O2 = CO2


nO2(needed) = (8.18 "\times" 3 + 12.27 "\times" 6.5 + 6.13 "\times" 0.5) - 0.82 = 106.54 (mol)

VO2(needed) = 106.54"\times" 22.4 = 2386.5 (L)

Vair(needed) = 2386.5 / 0.21 = 11364.3 (L) = 11.36 m3


Answer: Vair(needed) = 11.36 m3.